Show by giving a proof by contrapositive, that if 3n+2 is odd, then n is odd


1
Expert's answer
2021-10-27T09:49:25-0400

The contrapositive of the above statement is as follows.

If n isn't odd then 3n+2 isn't odd (If n is even then 3n+2 is even).

Let us prove the obtained statement:

If n is even then "n = 2k,\,k \in Z" . Then

"3n + 2 = 3 \cdot 2k + 2 = 6k + 2"

6k is even, 2 is even, so 6k+2 is even (as the sum of two even numbers).

Q. E. D.